Gamma rays have a frequency of 13 x 10^18 Hz. Calculate the wavelength of this gamma ray. 


1
Expert's answer
2021-03-18T14:03:52-0400
"\lambda=\dfrac{c}{f}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{13\cdot10^{18}\ Hz}=2.31\cdot10^{-11}\ m."